OpenStudy (anonymous):

find d4y/dx4,if y=log(2x+5)

myininaya (myininaya):

we need to find the 4th derivative

myininaya (myininaya):

have you found the first ?

myininaya (myininaya):

And I guess you log(2x+5) is understood to be log_(10)(2x+5)

OpenStudy (anonymous):

the first will be 2(2x+5)^-1

OpenStudy (anonymous):

is that it?

myininaya (myininaya):

Is the base of the log e?

myininaya (myininaya):

if so yes

OpenStudy (anonymous):

do i have to be finding the derivative one after the other?

myininaya (myininaya):

yep (Also is the base of that log e?)

myininaya (myininaya):

or is that log natural?

OpenStudy (anonymous):

is there no formula to find all the four at once?

myininaya (myininaya):

sometimes you are asked to derive a formula we can derive one here by observing pattern

OpenStudy (anonymous):

natural.

myininaya (myininaya):

Ok so you have f'=2(2x+5)^-1 now we need f''=

OpenStudy (anonymous):

f"=-1x2x2(2x+5)^-2

OpenStudy (anonymous):

is it?

myininaya (myininaya):

what is all those x's?

myininaya (myininaya):

oh are some of them multiplication signs and the only variable x you have is in that parenthesis next to the 2?

OpenStudy (anonymous):

multiplication

myininaya (myininaya):

You actually can see a pattern here and skip finding the third derivative

myininaya (myininaya):

notice: \[f'=2(2x+5)^{-1} ; f''=-1 \cdot 2^2 (2x+5)^{-2}\]

myininaya (myininaya):

let's take a guess at the fourth derivative from this

myininaya (myininaya):

you know the the derivatives are sign alternating

myininaya (myininaya):

I'm going to rewrite the first derivative one more time so it is more obvious \[f'=2^1(2x+5)^{-1} ; f''=-1 \cdot 2^2 (2x+5)^{-2} \]

myininaya (myininaya):

do you see a pattern yet?

myininaya (myininaya):

wait let me rewrite one more time to make more obvious \[f^{(1)}=2^1(2x+5)^{-1} ; f^{(2)}=-1 \cdot 2^2 (2x+5)^{-2}\]

OpenStudy (anonymous):

f'"=-1.-2.2^3(2x+5)^-3

myininaya (myininaya):

the 3rd derivative should be negative and shouldn't have 2*2^3 but just 2^3

myininaya (myininaya):

shouldn't be negative*

myininaya (myininaya):

oh wait is that the 4th derivative so many '

myininaya (myininaya):

that still wouldn't be the 4th derivative

OpenStudy (anonymous):

no, 3rd

myininaya (myininaya):

so you weren't able to see the pattern from the first two derivatives \[f^{(1)}=(-1)^{-1+1} \cdot 2^{1}(2x+5)^{-1}\] \[f^{(2)}=(-1)^{-1+2}\cdot 2^2 (2x+5)^{-2}\] \[f^{(3)}=(-1)^{-1+3} \cdot 2^3(2x+5)^{-3}\] now do you see the pattern?

myininaya (myininaya):

the (-1) to some power was thrown in there to get the alternating part

myininaya (myininaya):

notice the things that change and the things that don't change

myininaya (myininaya):

omg i made an error

myininaya (myininaya):

forgot to bring down the 2 I think

myininaya (myininaya):

\[f^{(3)}=(-1)^{-1+3} \cdot 2^{4}(2x+5)^{-3}\]

myininaya (myininaya):

so you were right i'm sorry

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

have got the pattern

OpenStudy (anonymous):

thank you

myininaya (myininaya):

you do so you could find the nth derivative? :)

myininaya (myininaya):

spoiler below ____for nth derivative: \[f^{(n)}=(-1)^{-1+n} \cdot 2^n \cdot (n-1)! \cdot (2x+5)^{-n}\]

myininaya (myininaya):

There is your formula if you really wanted to just find one for this problem.

myininaya (myininaya):

Also honestly there are too many unnecessary formulas to memorize like this one

myininaya (myininaya):

I say it is unnecessary because this can easily be derived by other formulas

myininaya (myininaya):

you can't remember them all our brain is limited

OpenStudy (anonymous):

yes

myininaya (myininaya):

You did very well @dzie so you get a medal too

OpenStudy (anonymous):

thank you myininaya